# Shoot from the HIP: Hessian Interatomic Potentials Without Derivatives

### v1

@@ A.4 Training Details
We train our model HIP-EquiformerV2 for roughly six days on only a single H100 GPU. The hyperparameters are listed in 3. We inherit the model settings and large parts of the optimizer settings from the HORM codebase (Cui et al., 2025). We expect much better results to be possible by training end-to-end, for longer, and tuning the hyperparameters.
Table 3: Hyperparameters used
@@
Trainer Epochs: 300 with limited batches (approx. 35 full epochs)
Limit train batches per epoch: 1600
A.5 Loss function
We are trying to design a loss function that emphasizes the lowest lying eigenvalues/eigenvectors. To design such a loss function, one could naively compute a loss directly as $L=\sum_{i}^{k}|\mathbf{v}_{i}^{\text{predict}}-\mathbf{v}_{1}^{\text{true}}|$. This requires computing the eigenvalues and eigenvectors of the true and predicted Hessian via an eigenspectrum decomposition. Unfortunately, backpropagation through such an eigenspectrum decomposition is numerically unstable (van der Aa et al., 2007). First, gradients computed using the eigenvectors tensor will only be finite when $\mathbf{H}$ has distinct eigenvalues. Furthermore, if the distance between any two eigenvalues is close to zero, the gradient will be highly sensitive, as it depends on the eigenvalues $\lambda_{i}$ through the computation of $\min_{i\neq j}\frac{1}{\lambda_{i}-\lambda_{j}}$. Instead, we propose the following subspace loss.
@@ A.5 Loss function
$\mathcal{L}_{\text{MAE+Sub}}$ $=\mathcal{L}_{\text{MAE}}+\mathcal{L}_{\text{sub}}^{(k=8,)}$ (36)
@@ A.6.1 Relaxations
We compare our RFO-based optimization (Besalú & Bofill, 1998) using predicted Hessians (RFO predicted), and our RFO with BFGS updates with initial predicted Hessian (RFO-BFGS predicted init) against first-order methods, quasi-second-order BFGS variants, and exact second-order approaches obtained via automatic differentiation or finite differences.
Our baselines include steepest descent with backtracking line search (SteepestDescent) and the first-order method FIRE (Bitzek et al., 2006). For quasi-second-order methods, we consider RFO with BFGS updates initialized from either the identity (RFO-BFGS unit init), AD Hessian (RFO-BFGS AD init), or finite difference Hessian (RFO-BFGS FiniteDifference init). For full second-order methods, we include RFO with either automatic differentiation or finite difference Hessians. We run relaxations on 80 reactant geometries from the Transition1x validation set (which is different from HORM-T1x), that we noise with $0.5$ Å RMS. We use the RS-RFO and redundant coordinate implementation in pysisyphus (Steinmetzer et al., 2021; Herbol et al., 2017). We set the convergence criteria to "Gaussian default" (see A.6.4), with a budget of 150 steps.
A.6.2 Transtion State Search with ReactBench
We use the recently introduced ReactBench benchmark (Zhao et al., 2025). The workflow consists of generating an initial guess using the growing string method (Aldaz et al., 2025; Zimmerman, 2015), followed by local search using the restricted-step partitioned rational-function-optimization method (RS-P-RFO) (Besalú & Bofill, 1998). Finally, convergence to the correct transition state is confirmed by frequency analysis, and following the intrinsic reaction coordinate (IRC) (Steinmetzer et al., 2021).
Following previous work (Zhao et al., 2025; Cui et al., 2025), we report success metrics of each step:
@@ A.6.2 Transtion State Search with ReactBench
    After RS-P-RFO, the frequency analysis determines geometry as a true transition state ("TS Success")
@@
    IRC converged to the same criteria and yields the original initial reactant and product ("IRC Verified").
@@ A.6.4 Convergence Criteria
We briefly describe the process of removing the redundant degrees of freedom of the Hessian (Wilson et al., 1955; Louck & Galbraith, 1976).
@@
The intrinsic reaction coordinate (IRC) traces the steepest-descent path on the potential energy surface (PES) from a transition state (TS) toward reactant and product minima. We use the IRC provided in the ReactBench benchmark (Zhao et al., 2025), which itself uses the pysisyphus implementation (Steinmetzer et al., 2021). We integrate in mass-weighted Cartesian coordinates to obtain a physically meaningful path, solving:
@@
where $\mathbf{x}$ are mass-weighted coordinates and $\mathbf{g}$ is the mass-weighted gradient. The pysisyphus implementation follows an Euler predictor-corrector (EulerPC) scheme (Hratchian et al., 2010). Starting from a TS, we diagonalize the mass-weighted Hessian, identify the imaginary mode (transition vector), and displace slightly along it to initialize the path. We perform small Euler integrations along $\mathbf{g}/\lVert\mathbf{g}\rVert$, updating the gradient via a local Taylor expansion using the current Hessian. The Hessian is updated with BFGS updates from $\Delta\mathbf{x},\Delta\mathbf{g}$. We refine the predicted point using distance-weighted interpolation (DWI) of previously visited points (coordinates, energies, gradients, Hessians). The corrector integrates the IRC equation on this surrogate PES using a modified Bulirsch-Stoer procedure with Richardson extrapolation.

### v2

@@ A.3 Batched AD Hessian performance
Figure 7: HIP compared to ground-truth DFT Hessian. The horizontal and vertical axis index the $(3N_{Atoms})^{2}$ entries of the Hessian. We depict the absolute values of test sample C5H6O in eV/Å2.
@@ A.4 Training Details
The hyperparameters our HIP-EquiformerV2 model are listed in 3. We inherit the model settings and large parts of the optimizer settings from the HORM codebase (cui2025horm).
@@
Trainer Steps: 500k (3-5 days on a H100 GPU)
Table 3: Model and training Hyperparameters. For simplicity, we use the same hyperparameters for HIP as HORM (cui2025horm).
Training resources
Training AD Hessians is exceptionally more expensive than training HIP. The memory cost of computing loss gradients through the full $3N\times 3N$ Hessian is prohibitive. Instead, the baseline models from cui2025horm randomly sample one or two columns of the Hessian via random Hessian-vector products at each training step. Despite this, training AD Hessians remains exceedingly expensive. The HORM paper states that they use 2 HVPs at a batch size of 128. Their checkpoint metadata shows that the per device batch size was 8, implying a 16xA30 GPU setup for 1 million gradient steps. In contrast, we trained HIP on a single H100 GPU for 500k steps. Additionally, each train step of HIP costs 1/5 the cost of AD Equiformer training on our hardware. Finally, we observe HIP fully converges after 300-500k training steps, while AD Hessians require a multiple of steps, due to the limited supervision of not being able to use the full Hessian.
Model Hessian Columns Batch Size BZ / GPU GPU # GPUs Training Steps
AlphaNet 1 32 16 H20 2 688,000
LeftNet-CF 1 64 16 A30 8 192,000
LeftNet-DF 2 64 16 A30 8 120,000
EquiformerV2 (E-F) 0 128 8 A30 16 384,000
EquiformerV2 2 128 8 A30 16 1,008,000
HIP-EquiformerV2 3N (all) 128 128 H100 16 500,000
Table 4: Summary of training resources. Comparing training with HIP to AD-Hessians with HIP as done by HORM (cui2025horm).
A.5 Additional results
A.5.1 Loss function
We are trying to design a loss function that emphasizes the lowest lying eigenvalues/eigenvectors. To design such a loss function, one could naively compute a loss directly as $L=\sum_{i}^{k}|\mathbf{v}_{i}^{\text{predict}}-\mathbf{v}_{1}^{\text{true}}|$. This requires computing the eigenvalues and eigenvectors of the true and predicted Hessian via an eigenspectrum decomposition. Unfortunately, backpropagation through such an eigenspectrum decomposition is numerically unstable (aaComputationEigenvalueEigenvector2007). First, gradients computed using the eigenvectors tensor will only be finite when $\mathbf{H}$ has distinct eigenvalues. Furthermore, if the distance between any two eigenvalues is close to zero, the gradient will be highly sensitive, as it depends on the eigenvalues $\lambda_{i}$ through the computation of $\min_{i\neq j}\frac{1}{\lambda_{i}-\lambda_{j}}$. Instead, we propose the following subspace loss.
@@ A.5.1 Loss function
$\mathcal{L}_{\text{MAE+Sub}}$ $=\mathcal{L}_{\text{MAE}}+\mathcal{L}_{\text{sub}}^{(k=8)}$ (44)
@@
RGD1
To test the extrapolation of HIP we test the same model checkpoints on the slightly larger molecules in the RGD1 dataset in table 9. Similar to T1x, models trained without Hessian data fail to produce accurate AD Hessians (EquiformerV2 (E-F)). HIP produces significantly better Hessians faster on every metric compared to AD. End-to-end training HIP improves over only training the readout head. We reach the same conclusions on the QM9 Hessian dataset in A.5.1. Training HIP from scratch is also orders of magnitude cheaper to train than AD Hessians, as we discuss in A.4.
eV/Å2 eV/Å2 unitless eV/Å2 ms
AD AlphaNet 0.456 4.008 0.734 0.545 935.4
LEFTNet-CF 0.243 1.388 0.821 0.337 1358.9
LEFTNet-DF 0.292 1.319 0.809 0.301 434.3
EquiformerV2 0.092 0.315 0.822 0.220 946.1
EquiformerV2 (E-F) 1.234 10.349 0.397 0.969 946.7
Predicted HIP-EquiformerV2 0.057 0.158 0.943 0.130 50.9
HIP-EquiformerV2 (end-to-end) 0.049 0.123 0.950 0.098 40.6
Table 9: HORM-RGD1 Comparison of Hessian prediction errors on the HORM-RGD1 validation set.
Computational cost compared to finite difference
Figure 9: Computational cost. Comparing (a) time and (b) memory of a forward without Hessians, with HIP, automatic differentiation (AD), finite difference Hessians. The memory for Forward Pass, FD Hessians, and HIP Hessians is nearly identical.
wander2025accessing showed that finite difference Hessians obtained from finetuned MLIPs can increase success rates in transition state search for catalysis. The advantage of finite difference is that it does not require additional memory over a regular forward pass, and only force labels are needed for training. The limitation of finite difference Hessians is their steep time cost. Like AD Hessians, finite difference requires $O(N^{2})$ time, but with an even larger prefactor, as can be seen in figure 9.
QM9 Hessian
[…]
Like the baseline, all models are trained end-to-end for each solvent separately. We train on one L40s GPU for one day for HIP and three days for AD Hessian variants. The training setup is listed in table 10. The results in table 11 show that HIP improves over all baselines, in particular the Nequip-based model from williams2025. EquiformerV2 models trained with AD Hessians via HVPs perform poorly even after three times the training wall time of HIP. Training without Hessians (E-F) yields poor accuracy on Hessian metrics.
Model Hessian Columns Batch Size Train Time / Epoch (s)
EquiformerV2 (E-F) 0 64 380
EquiformerV2 2 16 1900
EquiformerV2-CF 1 12 3200
HIP-EquiformerV2 3N 64 550
Table 10: Training configuration and epoch time for QM9 Hessian. Training was performed on a L40s GPU. Batch size is the maximum that could be fitted on the GPU.
Solvent Model Hessian Forces $\downarrow$ Hessian $\downarrow$ Eigenvalues $\downarrow$ CosSim ${\bm{v}}_{1}$ $\uparrow$ $\lambda_{1}$ $\downarrow$
eV/Å eV/Å2 eV/Å2 unitless eV/Å2
Vacuum E3x-Nequip AD 0.035 0.070 - - -
EquiformerV2 AD 0.030 0.340 2.641 0.425 0.151
EquiformerV2 (E-F) AD 0.002 2.107 32.471 0.076 0.175
EquiformerV2-CF AD 0.049 0.881 12.212 0.074 0.652
HIP-EquiformerV2 (end-to-end) Predicted 0.002 0.040 0.172 0.975 0.010
THF E3x-Nequip AD 0.028 0.067 - - -
EquiformerV2 AD 0.030 0.281 1.874 0.380 0.117
EquiformerV2 (E-F) AD 0.002 2.077 31.566 0.067 0.169
EquiformerV2-CF AD 0.048 0.846 10.447 0.079 0.523
HIP-EquiformerV2 (end-to-end) Predicted 0.002 0.060 0.240 0.878 0.033
Toluene E3x-Nequip AD 0.029 0.062 - - -
EquiformerV2 AD 0.030 0.320 2.232 0.424 0.104
EquiformerV2 (E-F) AD 0.002 2.095 31.606 0.072 0.158
EquiformerV2-CF AD 0.361 4.794 39.822 0.023 320.933
HIP-EquiformerV2 (end-to-end) Predicted 0.002 0.038 0.117 0.947 0.008
Water E3x-Nequip AD 0.028 0.054 - - -
EquiformerV2 AD 0.028 0.273 1.817 0.377 0.071
EquiformerV2 (E-F) AD 0.001 2.097 31.862 0.066 0.050
EquiformerV2-CF AD 0.068 0.991 11.440 0.070 1.290
HIP-EquiformerV2 (end-to-end) Predicted 0.001 0.025 0.091 0.926 0.012
Table 11: Comparison of Hessian prediction errors on the QM9 Hessian dataset. Denoted are the mean absolute error (MAE), cosine similarity (CosSim), and average time per forward pass on a L40s GPU. $\lambda_{1}$ and ${\bm{v}}_{1}$ where obtained after Eckart-projection. Bold values highlight the best-performing model. E3x-Neuqip models are taken from williams2025.
@@ A.6.1 Relaxations
We compare our RFO-based optimization (RSRFO1998a) using predicted Hessians (RFO predicted), and our RFO with BFGS updates with initial predicted Hessian (RFO-BFGS predicted init) against first-order methods, quasi-second-order BFGS variants, and exact second-order approaches obtained via automatic differentiation or finite differences.
Our baselines include steepest descent with backtracking line search (SteepestDescent) and the first-order method FIRE (bitzek2006). For quasi-second-order methods, we consider RFO with BFGS updates initialized from either the identity (RFO-BFGS unit init), AD Hessian (RFO-BFGS AD init), or finite difference Hessian (RFO-BFGS FiniteDifference init). For full second-order methods, we include RFO with either automatic differentiation or finite difference Hessians. We run relaxations on 80 reactant geometries from the Transition1x validation set (which is different from HORM-T1x), that we noise with $0.5$ Å RMS. We use the RS-RFO and redundant coordinate implementation in pysisyphus (steinmetzer2021; herbolComputationalImplementationNudged2017). We set the convergence criteria to "Gaussian default" (see A.6.4), with a budget of 150 steps.
A.6.2 Transition State Search with ReactBench
We use the recently introduced ReactBench benchmark (zhao2025). The workflow consists of generating an initial guess using the growing string method (aldazZimmermanGroupPyGSM2025; zimmerman2015), followed by local search using the restricted-step partitioned rational-function-optimization method (RS-P-RFO) (RSRFO1998a). Finally, convergence to the correct transition state is confirmed by frequency analysis, and following the intrinsic reaction coordinate (IRC) (steinmetzer2021).
Following previous work (zhao2025; cui2025horm), we report success metrics of each step:
@@ A.6.2 Transition State Search with ReactBench
    After RS-P-RFO, the frequency analysis determines geometry as a true transition state, i.e. exactly one negative eigenvalue of the Hessian ("TS Success")
@@
    Both IRC calculations converged and land at the original initial reactant and product that were used as the starting points for GSM ("IRC Verified").
@@ A.6.4 Convergence Criteria
We briefly describe the process of removing the redundant degrees of freedom of the Hessian (wilson1955; louck1976).
@@
The intrinsic reaction coordinate (IRC) traces the steepest-descent path on the potential energy surface (PES) from a transition state (TS) toward reactant and product minima. We use the IRC provided in the ReactBench benchmark (zhao2025), which itself uses the pysisyphus implementation (steinmetzer2021). We integrate in mass-weighted Cartesian coordinates to obtain a physically meaningful path, solving:
@@
where $\mathbf{x}$ are mass-weighted coordinates and $\mathbf{g}$ is the mass-weighted gradient. The pysisyphus implementation follows an Euler predictor-corrector (EulerPC) scheme (hratchianSteepestDescentReaction2010). Starting from a TS, we diagonalize the mass-weighted Hessian, identify the imaginary mode (transition vector), and displace slightly along it to initialize the path. We perform small Euler integrations along $\mathbf{g}/\lVert\mathbf{g}\rVert$, updating the gradient via a local Taylor expansion using the current Hessian. The Hessian is updated with BFGS updates from $\Delta\mathbf{x},\Delta\mathbf{g}$. We refine the predicted point using distance-weighted interpolation (DWI) of previously visited points (coordinates, energies, gradients, Hessians). The corrector integrates the IRC equation on this surrogate PES using a modified Bulirsch-Stoer procedure with Richardson extrapolation.